OpenStudy (anonymous):

Simplify: x^11/3 / x^2/3 (Can't remember to save my life haha, easy medal)

OpenStudy (anonymous):

I know it equals x^3 when simplified, I just can't remember the steps.

OpenStudy (***[isuru]***):

\[\frac{ a^{b} }{ a^{c} } = a^{b - c}\]

OpenStudy (anonymous):

Okay, so 11/3 - 2/3 = 9/3 = 3?

OpenStudy (***[isuru]***):

\[\frac{ x^{\frac{ 11 }{ 3 }} }{ x ^{\frac{ 2 }{ 3 }} } = x^{\frac{ 11 }{ 3 } - \frac{ 2 }{ 3 }}\]

OpenStudy (***[isuru]***):

yeah... that how the things happens :)

OpenStudy (anonymous):

Thank you! Would mind helping me with my last one?

OpenStudy (anonymous):

@***[ISURU]***